### v1

# SPGD: Steepest Perturbed Gradient Descent Optimization

Abstract
Optimization algorithms are pivotal in advancing various scientific and industrial fields but often encounter obstacles such as trapping in local minima, saddle points, and plateaus (flat regions), which makes the convergence to reasonable or near-optimal solutions particularly challenging.
This paper presents the Steepest Perturbed Gradient Descent (SPGD), a novel algorithm that innovatively combines the principles of the gradient descent method with periodic uniform perturbation sampling to effectively circumvent these impediments and lead to better solutions whenever possible. SPGD is distinctively designed to generate a set of candidate solutions and select the one exhibiting the steepest loss difference relative to the current solution. It enhances the traditional gradient descent approach by integrating a strategic exploration mechanism that significantly increases the likelihood of escaping sub-optimal local minima and navigating complex optimization landscapes effectively. Our approach not only retains the directed efficiency of gradient descent but also leverages the exploratory benefits of stochastic perturbations, thus enabling a more comprehensive search for global optima across diverse problem spaces. We demonstrate the efficacy of SPGD in solving the 3D component packing problem, an NP-hard challenge. Preliminary results show a substantial improvement over four established methods, particularly on response surfaces with complex topographies and in multidimensional non-convex continuous optimization problems. Comparative analyses with established 2D benchmark functions highlight SPGD’s superior performance, showcasing its ability to navigate complex optimization landscapes. These results emphasize SPGD’s potential as a versatile tool for a wide range of optimization problems.
1 Introduction
Mathematical optimization is a fundamental process in engineering, science, and economics. Its main objective is to find solutions that minimize a predefined objective, typically expressed in terms of a real-valued function, while adhering to given constraints. This pursuit of optimal solutions is crucial in solving complex problems, where achieving the best possible results necessitates a careful balance of numerous factors and variables.
Among the many optimization techniques available, the gradient descent (GD) method stands out as a foundational and extensively used tool, and its origins can be traced back to Cauchy’s pioneering work [1]. However, despite its widespread use, the gradient descent method has certain limitations. One of its major drawbacks is its tendency to get trapped in sub-optimal states, including saddle points and local minima, which may offer minimal improvement in solution quality. Additionally, the method may encounter difficulties in making progress towards the desired outcome when faced with flat regions in the problem space.
To address these challenges, extensive research efforts have been focused on enhancing the performance of the gradient descent method. As a result, numerous variants have been developed, each specifically designed to overcome the aforementioned pitfalls [2]. One notable variant is the Perturbed Gradient Descent (PGD), which has gained attention for its ability to navigate away from saddle points and potentially converge towards second-order optimal points [3].
In this paper, we present a strategic randomized perturbation algorithm combined with the gradient descent method, leveraging the strengths of both exploring the search space through randomized perturbation and converging to optimal points using gradient information. By introducing cyclical perturbations, our approach strategically balances the need for exploration with the efficiency of exploitation. Moreover, applying perturbations periodically rather than at every iteration significantly reduces computational costs, making the optimization process more efficient without sacrificing the thoroughness of the search. It promises a more reliable pathway to discovering superior solutions, thereby expanding the horizon of possibilities in optimization challenges. This enhanced method is designed not only to navigate more effectively through the complexities of practical optimization landscapes but also to refine the search for optimal solutions with greater precision.
The remainder of this paper systematically explores the Steepest Perturbed Gradient Descent (SPGD) algorithm and its comparative advantages in the domain of optimization. Section 2 delves into a variety of related methodologies, focusing on variants of the gradient descent method and the integration of perturbation sampling techniques. These approaches establish a foundation for understanding the landscape of optimization strategies and highlight the necessity for innovations, such as SPGD. Section 3 is dedicated to a detailed exposition of the SPGD algorithm itself, including its theoretical underpinnings, algorithmic structure, and the rationale behind its design choices. Following this, Section 4 presents numerical results from a series of experiments designed to evaluate the performance of SPGD against various established optimization algorithms. These experiments are conducted on a selection of well-known optimization test functions, providing a rigorous comparison and demonstrating the practical implications of SPGD in addressing complex optimization challenges. Finally, Section 5 discusses the outcomes of these comparisons, emphasizing the superior performance of SPGD over the methods analyzed. The conclusions not only underscore the effectiveness of SPGD but also set the stage for future research directions and potential applications in broader optimization contexts.
2 Related Work
Simulated annealing (SA) [4] and genetic algorithm [5] are heuristic optimization algorithms that use randomness and selection mechanisms inspired by natural processes to explore the solution space and select the best candidates for further iteration. These algorithms can be used for different types of optimization problems such as continuous, discrete, non-convex, and multi-objective problems [6, 7]. These methods may require significant computational resources and careful tuning of parameters (e.g., temperature in simulated annealing or mutation rate in evolutionary algorithms) to balance exploration and exploitation effectively.
The gradient descent (GD) method is a first-order optimization algorithm that updates the design variables in the direction opposite to the gradient of the objective function with respect to those variables [2]. It’s widely used due to its simplicity and efficiency in convex problems. The gradient descent method converges to a local optimal solution with a mathematical guarantee. Gradient descent tends to exploit local information to improve the solution iteratively. However, it may not explore the search space effectively, potentially getting trapped in local minima or saddle points, particularly in non-convex optimization landscapes. It struggles with flat areas where the gradient is close to zero, leading to slow or no progress [3].
Nesterov’s Accelerated Gradient (NAG) method enhances traditional gradient descent by incorporating a forward-looking step. This tweak allows the optimizer to anticipate future gradients, reducing oscillations and speeding up convergence, particularly in convex settings. NAG is highly effective in training deep neural networks due to its efficiency in navigating high-dimensional data spaces. However, its performance can vary in non-convex environments with complex landscapes [8, 9].
In the exploration of hybrid optimization methods, a notable approach combines the exploratory strengths of Simulated Annealing (SA) with the precise, local search capabilities of Gradient Descent (GD). This method strategically employs SA to break free from local optima by conducting a thorough search for a more promising solution candidate, upon which GD resumes. While this synergy offers a dynamic pathway to escape local minima, it introduces a significant computational burden. Moreover, this method diverges from traditional GD in that it cannot rely on the norm of the gradient as a criterion for termination. This alteration results in a less stringent stop condition, potentially affecting the algorithm’s efficiency and termination reliability [10].
Another hybrid technique is perturbed gradient descent (PGD) that addresses the challenge of stagnation—a state where the gradient becomes negligible, and no further progress seems attainable in optimizing the objective function. This method introduces a single perturbation to the current solution when progress halts, effectively nudging the search process out of stagnation before proceeding with GD. This approach demonstrates an ability to escape saddle points effectively [3, 11, 9]. However, its performance is notably diminished in flat regions of the search space, where such perturbations fail to provide a meaningful direction for improvement.
In addition to the deterministic and heuristic methods previously discussed, the random walk method offers a stochastic approach to optimization that is particularly advantageous in complex, non-convex landscapes. Random walks operate by making a sequence of moves, each determined randomly in terms of direction and step size. This method inherently avoids the common pitfalls of gradient-based approaches, such as becoming trapped in local minima, by facilitating an unbiased exploration of the solution space. This characteristic is critical when dealing with high-dimensional optimization problems where the landscape is riddled with numerous local optima and saddle points [12]. Despite their potential for encompassing space exploration, random walks are often criticized for their inefficiency and slow convergence, especially in large-scale problems. They require a large number of iterations to approach the vicinity of a global optimum, as their exploration process lacks directionality inherent to methods like gradient descent or even simulated annealing. To address these limitations, researchers have explored hybrid strategies that combine the exploratory strengths of random walks with more systematic search techniques to balance exploration with exploitation more effectively [13, 14].
3 SPGD
Traditional gradient descent algorithms efficiently exploit local gradient information to improve solutions iteratively. To minimize a given function , the updating rule at each iteration is [3]:
where is the number of current iteration, and is the step size, and is the gradient of .
However, in non-convex high-dimensional problems, the gradient descent method can become trapped in local minima or saddle points, missing out on globally optimal configurations. To address this limitation, we propose a novel algorithm that combines gradient descent with periodic randomized perturbations. These perturbations are particularly effective in non-convex, high-dimensional problems, where even small modifications can significantly alter the solution’s position within the search space. This sensitivity to perturbations is crucial in navigating the complex terrain of such problems, where the landscape of potential solutions is riddled with local optima. By introducing strategically randomized perturbations, our algorithm enhances its ability to escape these local optima, thereby facilitating a more extensive exploration of the solution space. This periodic application of perturbations is key to avoiding the oscillatory behavior often observed in optimization trajectories of sampling-based methods, which can lead to inefficiencies and slow convergence. This approach proves particularly advantageous in complex optimization scenarios characterized by challenges such as flatness, ruggedness, or saddle points of the objective surface, where conventional optimization algorithms might falter in making meaningful progress. These perturbations are drawn from uniform random distributions111Having a uniform distribution allows us to sample the solution space around the current solution uniformly within the range of . This approach leads to more explorative sampling by equally covering the vicinity around the current position, rather than concentrating on the immediate area around the current solution or extending far beyond it. with constant amplitude profiles222The main reason for choosing a constant amplitude profile is to maintain the simplicity of the algorithm in these benchmark functions. However, the amplitude profile () can be adjusted to any arbitrary profile based on the specific requirements of the optimization problem at hand.. The uniform random distribution will create perturbed candidates around the gradient descent solution every iterations. All perturbed candidates will be evaluated and compared with the gradient descent solution. If the minimum value of perturbed candidates is equal or less than the value of the gradient descent solution, the corresponding perturbed candidate will be selected as the new solution. This policy of accepting solutions with equal objective values intentionally increases the algorithm’s emphasis on exploration over exploitation within the optimization process. Such an approach is particularly advantageous in scenarios where the objective surface is flat, and traditional gradient descent methods stall due to insufficient gradient information. By facilitating exploration in these flat regions, SPGD ensures continued progress towards finding a global optimum, preventing the algorithm from becoming prematurely anchored to suboptimal solutions. The pseudo code of SPGD algorithm is described in Algorithm 1.
In the SPGD algorithm, the method of applying perturbations is adaptable to the specific requirements of the optimization problem at hand. For unconstrained optimization problems, such as 2D test function benchmarks and neural network training, perturbations are applied simultaneously to all variables, utilizing a uniform distribution with a constant amplitude. This ensures a broad, uniform exploration of the solution space, which is generally suitable for the landscapes presented by these types of problems.
However, when dealing with constrained problems like the 3D component packing, which present a complex optimization landscape, a different approach is warranted. In such scenarios, perturbation of a single variable can potentially lead to an infeasible solution or a worse candidate due to the constraints involved. To mitigate this, each variable is perturbed separately, effectively reducing the complexity and dimensionality of the optimization problem by focusing on one variable at a time, with all others held constant [15]. This targeted perturbation allows for a more controlled exploration of the solution space, ensuring that the search remains within feasible regions and is more likely to improve upon the current solution. This adaptive feature is designed to tailor the exploration process more precisely to the problem’s landscape, enhancing the algorithm’s flexibility and effectiveness in navigating constrained environments.
The parameters of SPGD, notably the number of perturbations , the perturbation interval , and the perturbation amplitude , play crucial roles in shaping the algorithm’s behavior and performance. Increasing enhances the likelihood of discovering superior solutions by broadening the search during perturbation phases, albeit at a higher computational cost. A smaller amplifies the algorithm’s exploratory behavior, contributing to a more thorough search of the solution space but also increasing computational demands and leading to more oscillatory convergence patterns. Conversely, selecting a larger facilitates wider exploration of the search space, though its effectiveness is highly contingent on the specific problem being addressed. For problems where small variations in inputs lead to significant changes in outputs, a large amplitude may not yield beneficial results, underscoring the importance of parameter tuning to align with the problem’s characteristics.
To demonstrate the execution flow of the SPGD algorithm and to provide intuitive insights into its operation, Figure 1 offers a detailed visualization of a simple 1D parabolic landscape. This example illustrates the iterative process where traditional gradient descent, represented by the green line, begins the optimization journey. Upon activation of the perturbation process after iterations, SPGD generates candidates (depicted as blue pluses) around the current position. These candidates explore the vicinity of the solution space, identifying potential points with lower function values. The most promising candidate, offering the lowest function value, is then compared to the current position. As depicted, this candidate yields a better function value and is thus chosen, causing the algorithm to make a significant ”jump” to this new point. The process then resumes with gradient descent steps leading to faster convergence to the optimal solution. This example effectively showcases how SPGD combines exploration and exploitation to enhance convergence speed in scenarios where traditional gradient descent may progress slowly.
4 Numerical Results
We present here a thorough evaluation of the proposed Steepest Perturbed Gradient Descent (SPGD) algorithm, comparing its performance against several established optimization methods. The comparison includes traditional gradient descent (GD), Perturbed Gradient Descent (PGD), MATLAB function which is a versatile solver for constrained optimization problems [16], and Simulated Annealing (SA) [17].
Our initial analysis is conducted through the lens of four challenging 2D benchmark functions, selected for their known difficulties and relevance in assessing optimization algorithms’ efficacy. These test functions are recognized benchmarks within the optimization community, providing a diverse set of landscapes to evaluate each algorithm’s ability to navigate complex, non-convex, and potentially deceptive optimization spaces [18]. For each test function, we apply , Simulated Annealing, traditional gradient descent, PGD, and SPGD, meticulously recording and analyzing the results. In both the SPGD and PGD algorithms, the amplitude of the perturbation is set to be the same, ensuring comparability, is set to , and is set to . Furthermore, the gradient function is explicitly passed to the function in each scenario to guide the optimization process.
Key performance indicators include the accuracy of the solution, measured by the proximity to the known global optimum[19]; the computational efficiency, quantified by the number of function evaluations and CPU execution time. The SPGD algorithm’s source code, along with comparative analyses against the discussed methods using additional 2D challenging test functions, is publicly accessible on GitHub 333Source code and comparisons available at: https://github.com/Amir-M-Vahedi/SPGD-Benchmark-Functions.
Test function 1
The MATLAB Peaks function [20] presents a formidable challenge for optimization algorithms due to its intricate landscape, which features one global minimum, multiple local minima, a saddle point, and extensive flat regions. This complexity makes the Peaks function a critical benchmark for assessing the capabilities of optimization techniques, particularly those based on gradient descent. Traditional gradient descent methods often struggle with such landscapes, as they can easily become trapped in local minima or stall in flat areas, failing to make significant progress towards the global optimum [21]. The mathematical expression defining the Peak test function is given as follows:
It has a global minimum point located at , with an optimal function value of . The initial condition is chosen randomly to be , and the is set to 2.5. Figure 2 and 3 illustrate the 3D view and top view of optimization trajectory across the Peaks function surface. The convergence history comparison is shown in Fig 4. The total number of function evaluations, the converged optimal value, and CPU execution time for different methods are given in Table 1. Based on the results depicted in Figures 2, 3, and 4, and performance metrics in Table 1, it is evident that the GD and PGD algorithms become trapped in local minima. In contrast, the , SA, and SPGD algorithms successfully converge to the global optimum. Among these three, SPGD demonstrates the lowest computational cost. Notably, despite the method having fewer function evaluations, its CPU time is approximately 20 times greater than that of the SPGD algorithm.
| Algorithm | Total Fun. Evaluations | CPU Time[ms] | |
| GD | 1472 | -3.0498 | *6.85 |
| PGD | 1599 | -3.0498 | *6.37 |
| 60 | -6.5511 | 123.68 | |
| SA | 1341 | -6.5511 | 131.67 |
| SPGD | 274 | -6.5511 | 6.18 |
Test function 2
The Ackley function is a well-known non-convex optimization benchmark that poses a significant challenge to optimization algorithms, particularly due to its deceptive landscape characterized by a global optimum surrounded by a multitude of local minima [22]. This function is specifically designed to test the ability of optimization methods to escape local minima and efficiently search for the global optimum in a complex, multidimensional space. The Ackley function’s landscape features a large number of local minima leading towards the global minimum, making it an exemplary test case for evaluating the robustness and effectiveness of algorithms against the risk of premature convergence. The global minimum of the Ackley function is located at the origin (, ), with an optimal function value of zero (), which further serves as a clear target for optimization efforts. The formula representing the Ackley test function is articulated below:
The initial condition is chosen randomly to be , and the is set to 2.5. Figure 5 and 6 illustrate the 3D view and top view of optimization trajectory across the Ackley function surface. The convergence history comparison is shown in Fig 7. The total number of function evaluations, the converged optimal value, and CPU execution time for different methods are given in Table 2. Taking into account the data presented in the mentioned figures and table, analysis of the Ackley test function reveals that the GD, PGD, and methods became ensnared in local minima. In contrast, only the SA and SPGD algorithms successfully navigated to the global solution. Among these, SPGD not only achieved convergence with greater precision—approaching closer to the global optimum—but also demonstrated a computational speed, with CPU execution time being six times faster than its counterpart.
| Algorithm | Total Fun. Evaluations | CPU Time[ms] | |
| GD | 327 | 9.3933 | *6.38 |
| PGD | 477 | 9.3933 | *6.01 |
| 24 | 9.3530 | *107.23 | |
| SA | 504 | 3.32-4 | 59.99 |
| SPGD | 1501 | 8.10-5 | 10.31 |
Test function 3
The Easom function stands as a notable unimodal steep ridge [18] test function within the realm of optimization, particularly distinguished by its singular global optimum that resides in an extensive flat area. This flat region is characterized by minimal gradient variations, presenting a unique challenge for optimization algorithms, especially those reliant on gradient information to navigate the search space. The function is defined over a domain of (, ) for both and dimensions, emphasizing the necessity for optimization techniques to efficiently explore large search areas to locate the optimum [23]. The significance of the Easom function as a test scenario with simple mathematical formulation lies in its ability to simulate real-world optimization problems where the solution space is largely homogeneous, yet contains a singular, critical point of interest. This function tests the exploration strategies of algorithms, challenging them to avoid the pitfalls of vast non-informative regions. It emphasizes the importance of balance between exploration and exploitation, as effective optimization methods must not only navigate vast spaces efficiently but also recognize and converge to the global optimum with high precision. Mathematically, the Easom function’s global optimum is uniquely situated at (, ), where it attains a value of . The formula of the Easom test function is provided below:
The initial condition is chosen randomly to be , and the is set to 5. Figure 8 and 9 illustrate the 3D view and top view of optimization trajectory across the Easom function surface. The convergence history comparison is shown in Fig 10. The total number of function evaluations, the converged optimal value, and CPU execution time for different methods are given in Table 3. Reflecting on the performance metrics for the Easom test function, it is evident that only the SPGD algorithm successfully pinpointed the global optimum. As anticipated, GD was hindered in its progression by the minimal gradient values inherent to the function’s extensive flat regions. Similarly, both GD and the method failed to escape these flat expanses, effectively becoming ensnared within them. Among the competing methods, only SA managed to navigate towards a more favorable outcome, yet it fell short of achieving convergence to the global optimum, underscoring the distinctive effectiveness of SPGD in this scenario.
| Algorithm | Total Fun. Evaluations | CPU Time[ms] | |
| GD | 1 | 0 | *0.1 |
| PGD | 2021 | 0 | *5.18 |
| 1 | 0 | *51.15 | |
| SA | 1009 | -3.38-160 | *99.45 |
| SPGD | 6001 | -1 | 14.86 |
Test function 4
The Levy Function No. 13, characterized by its multimodality and non-convexity, presents a unique challenge for optimization algorithms with its single global optimum amidst a noisy, periodic distribution of local minima. This function tests an algorithm’s precision in distinguishing the global optimum from numerous suboptimal states, a key trait for solving complex real-world problems. It serves as a critical benchmark for evaluating the balance between exploration and exploitation in optimization techniques, underscoring its significance in both theoretical and practical applications. The global optimum of this function is strategically located at (, ), where it attains a value of . The expression for the Levy Function No. 13 is detailed below [24]:
The initial condition is chosen randomly to be , and the is set to 2.5. Figure 11 and 12 illustrate the 3D view and top view of optimization trajectory across the Levy Function No. 13 surface. The convergence history comparison is shown in Fig 13. The total number of function evaluations, the converged optimal value, and CPU execution time for different methods are given in Table 4. Based on the performance analysis for this test function, the GD, PGD, and methods were unable to find the global optimum, getting stuck in local minima instead. Notably, settled in a particularly poor local minimum. In contrast, both Simulated Annealing (SA) and SPGD successfully navigated to the global optimum. However, SPGD distinguished itself by achieving a more accurate solution, requiring fewer function evaluations, and demonstrating a 25 times faster CPU execution time compared to SA.
| Algorithm | Total Fun. Evaluations | CPU Time[ms] | |
| GD | 2001 | 6.2915 | *4.58 |
| PGD | 2001 | 6.2915 | *3.85 |
| 20 | 30.5009 | *88.90 | |
| SA | 2018 | 6.78-7 | 175.38 |
| SPGD | 1760 | 2.45-13 | 7.56 |
The challenges presented by these test functions, including their rugged landscapes and deceptive local minima, contain features that bear resemblance to those encountered in the energy landscape of protein folding. This complex biological process is characterized by a similarly intricate energy landscape that features multiple local optima (kinetic traps), rugged terrain, and steep energy barriers (sharp valleys and hills)[25, 26, 27, 28], as depicted in Figure 14.
The SPGD algorithm’s performance on these test functions suggests its potential utility in addressing the complex optimization problems inherent in protein folding. By adeptly navigating through challenging landscapes to find global or near-global optima, SPGD could significantly contribute to bioinformatics and molecular biology by optimizing protein structures to understand their function and interactions more accurately.
This analogy not only highlights the broader applicability of SPGD but also underscores the importance of developing robust optimization techniques that can effectively deal with the complexities of both mathematical functions and biological systems [29, 30].
Expanding our investigation beyond conventional 2D test functions, we also apply our algorithm (SPGD) to a 3D component packing problem, a task distinguished by its NP-hard classification [31, 32]. This problem introduces a unique set of challenges, including flat area saddle points and local optima, that further test the robustness of our approach against traditional gradient descent and simulated annealing methods.
3D Component Packing Problem
In the 3D component packing problem, we focus on arranging 3D objects with arbitrary shapes as compactly as possible without collision, akin to a simplified version of the interconnected systems with physical interactions (SPI2) problem but without considering the routing interconnections between objects [33]. Optimization methods often face challenges in this landscape, such as getting trapped in local minima or stalling in flat areas, thus failing to advance significantly towards the global optimum. The non-convex nature of the objective function, characterized by multiple local optima and saddle points, poses substantial challenges to any standard optimization technique. Nonetheless, our SPGD algorithm, which integrates randomized perturbations, is tailored to navigate these complex landscapes more effectively, demonstrating its adaptability and enhanced performance compared to conventional techniques.
Our 3D packing scenarios presented here are specialized instances of those tackled by SPI2-F – a novel and more general packing and layout optimization presented in [34], that performs both packing and layout optimization of complex interconnected systems in a multi-physics environment. The specialized scenarios presented below have been chosen to include cases that have known global optima. We note that an efficient packing method based on Fast Fourier Transforms444Fast Fourier Transforms have been previously shown to offer an elegant and efficient approach to compute collisions and penetrations as well as shape complementarity [35, 36, 37]. was introduced recently in [38], which restricts the orientation of the objects to an axis-alignment and hence allows rotations in 90-degree increments. By contrast, our method allows arbitrary rotations and alignments in space.
In the 3D component packing problem, our primary objectives are twofold: minimize the volume of the bounding box containing the components () while avoiding collisions between the components. Therefore, we define the mathematical expression of the objective function as follows:
where represents the weight associated with the bounding box volume, 1-4 is the weight related to collision avoidance, 1-5 is a small value to avoid singularity, and denotes the minimum distance between the spheres of different components.
The complexity and high dimensionality of this problem are underscored by the fact that each object in our example consists of spheres, and each component is controlled by six variables – three for displacement and three for orientation. The problem also incorporates constraints related to collision avoidance. To effectively navigate the highly non-convex and constrained space of the component packing problem, our approach involves tailored adaptations to the perturbation mechanism used in the Steepest Perturbed Gradient Descent (SPGD) algorithm. Perturbations are applied separately to the components’ displacement and orientation, ensuring a thorough optimization of both aspects of component placement.
In the early iterations, we enhance the exploration and facilitate the escape from suboptimal solutions by accepting solutions with worse volume outcomes by a prescribed factor. This acceptance factor decreases in a linear profile over the iterations until it reaches , at which point the algorithm only accepts new solutions that have the same or lower volume, thus refining the search towards the most compact configurations. Additionally, the amplitude of the perturbations for both displacement and orientation is controlled through a lower-bounded linear profile, which ensures that perturbations decrease in magnitude as the optimization process progresses, aligning more closely with the finer adjustments needed as the solution space is narrowed down. To further optimize the perturbation process and avoid ineffective perturbations, especially in cases where objects are too close to each other to allow for meaningful spatial adjustments, the frequency of perturbations is reduced using a lower bounded linear profile. This adaptive frequency adjustment helps prevent unnecessary computational expenditure on perturbations that are unlikely to be accepted due to collision constraints. These strategic adaptations enable SPGD to more effectively handle the complexities of packing diverse objects into a constrained space, making it robust against the challenges posed by the non-convex nature of the problem.
The implementation of this algorithm is carried out in Python using the PyTorch framework, which leverages CUDA for accelerated computation on GPUs. This setup allows for substantial improvements in computational efficiency, essential for managing the high-dimensional space of this problem. Using torch.autograd, we automatically compute the partial derivatives of the loss function with respect to the displacement and orientation vectors of each component. This gradient information is then used to update the component positions and orientations according to the update rule of gradient descent (1), akin to methods typically employed in deep learning optimizations. To further enhance the exploration capabilities of the optimization process, the sequence of component perturbations is shuffled in each iteration, promoting a more robust search through the solution space. In evaluating the effectiveness of the SPGD algorithm, we conducted a comparative analysis with the traditional Gradient Descent (GD) method across a series of increasingly complex packing scenarios.
The scenarios were designed to assess both algorithms under various conditions, ranging from uniform object sizes to irregular and diverse shapes, thereby testing their adaptability and efficiency in real-world packing challenges. Our implementation of Simulated Annealing diverged rather than converged, particularly in complex scenarios. This divergence can largely be attributed to the restrictive collision constraints integrated within the objective function (6), which prevent objects from moving through each other. Unlike the approach taken in reference [32], where collision constraints were relaxed and followed by refinement steps, our implementation maintained these constraints, leading to no evident signs of convergence and indicating the unsuitability of Simulated Annealing for these applications.
Moreover, Perturbed Gradient Descent (PGD) was not utilized in the 3D packing problem comparison. The reason for this is twofold: firstly, the norm of the partial derivative vector in this problem setting does not approach zero due to the direct inclusion of collision constraints within the objective function. Secondly, the primary cause for algorithm termination is often the occurrence of collisions between objects, which deviates from the typical operational premise of PGD. Additionally, PGD’s poor performance in separate 2D benchmark functions, which feature complex and challenging loss landscapes, further illustrates its limitations in navigating complicated optimization scenarios. This combination of factors reaffirms the decision to exclude PGD from the comparative analysis in our 3D packing problem.
4.1 Initial Configuration and Setup
Before delving into the comparative results, it is essential to note that both the SPGD and GD algorithms were initiated from the same configuration in each scenario to ensure a fair comparison. The initial setup involved distributing the objects well within the 3D space, providing sufficient free space around each object to avoid immediate collisions. Furthermore, the orientations of the objects were randomly chosen, introducing additional complexity and ensuring that the problem remained challenging for the optimizers. This initialization strategy was crucial for testing the algorithms’ abilities to effectively explore and optimize from a non-advantageous starting point.
4.2 Experimental Scenarios and Results
The following scenarios were considered for the comparison:
Scenario 1: Four identical cubes, where the global optimum is known, and cubes are packed together with the same orientation.
Scenario 2: Eight identical cubes, testing the scalability and spatial reasoning of the algorithms.
Scenario 3: Eight cubes of varying sizes, introducing a non-uniform configuration without a known global optimum, to evaluate heuristic capabilities.
Scenario 4: Eight objects of different complex shapes (gears, hooks, rivets, etc.), representing an industrial challenge with an unknown optimal packing configuration.
4.3 Analysis of Scenario 1: Four Identical Cubes
In Scenario 1, the initial configuration of the four identical cubes is depicted in Figure 15. This setup was designed to test each algorithm’s ability to navigate a relatively simple scenario where the global optimum involves aligning all cubes in a compact configuration. The results of the final configurations found by the GD and SPGD algorithms are illustrated in Figure 16, showing both Gradient Descent and Steepest Perturbed Gradient Descent results side by side.
The outcomes depicted in the figures reveal that, due to the collision constraint, GD struggled to converge to the global solution and settled in a suboptimal local minimum. In contrast, SPGD successfully converged to the global optimal configuration, effectively avoiding local minima and fulfilling the collision constraints more efficiently. To further illustrate the performance dynamics over the course of the optimization, the loss convergence history for both algorithms is plotted in Figure 17. This figure shows loss values as a function of elapsed time and the number of iterations.
Although SPGD achieved the optimal configuration more rapidly in terms of the number of iterations, it required more computational time overall compared to GD. These plots (Figures 17) help demonstrate that while SPGD’s iterations are more effective at progressing toward the global optimum, they are computationally more intensive, likely due to the complexity of the perturbation calculations and the more sophisticated collision checks involved.
This scenario underscores SPGD’s strengths in effectively navigating optimization landscapes with collision constraints and its ability to reach global optima where traditional GD may fail. However, the increased computational demand highlights an area for further optimization and efficiency improvements in SPGD’s implementation.
4.4 Analysis of Scenario 2: Eight Identical Cubes
In Scenario 2, the initial configuration of eight identical cubes is depicted in Figure 18. This scenario was designed to assess each algorithm’s ability to scale and manage increased numbers of objects while maintaining an efficient packing configuration. The outcomes of the final configurations found by the GD and SPGD algorithms are illustrated in Figures 19(a) and 19(b), respectively.
During the optimization process, the GD algorithm encountered significant issues and ceased further packing adjustments due to a collision between the yellow and red cubes, effectively stopping the optimization prematurely. In contrast, the SPGD algorithm managed to navigate around this problem and did not converge to the global optimal solution but found a notably more compact suboptimal solution, approximately three times more space-efficient than the configuration found by GD.
To further illustrate the performance dynamics over the course of the optimization, the loss convergence history for both algorithms is plotted in Figure 20, showing loss values as a function of elapsed time and the number of iterations.
Although SPGD did not achieve the global optimum, it provided a significant improvement over GD by finding a much more compact solution rapidly. This scenario demonstrates SPGD’s superior capability in effectively navigating complex landscapes and managing collision constraints dynamically compared to GD. The increased performance in finding a substantially better solution highlights the potential of SPGD for more effective space utilization in packing problems.
4.5 Analysis of Scenario 3: Eight Cubes of Different Sizes
In Scenario 3, which introduces a higher level of complexity due to the use of eight cubes of different sizes, the initial configuration is shown in Figure 21. This setup challenges the algorithms’ ability to efficiently manage and optimize space in a more heterogeneous environment.
The SPGD algorithm’s performance in this scenario was notably superior, as it converged to a more compact solution significantly faster than the traditional GD method. The results of the final configurations found by the GD and SPGD algorithms are shown in Figures 22(a) and 22(b), respectively.
Despite the lack of a known global optimal solution due to the varying sizes and potential configurations, SPGD effectively utilized its perturbation mechanism to explore and optimize the packing arrangement. This scenario highlights the algorithm’s adaptability and efficiency in handling diverse object dimensions, which is crucial for real-world applications.
To further evaluate the performance dynamics, the loss convergence history for both algorithms is plotted in Figure 23, showing loss values as a function of elapsed time, and the number of iterations.
These figures demonstrate that SPGD not only achieves a more desirable outcome but also does so with greater computational efficiency in terms of iteration count, despite the complex interplay of different-sized objects. This efficiency underscores SPGD’s potential as a robust tool for tackling sophisticated packing challenges where traditional methods might falter.
4.6 Analysis of Scenario 4: Eight Objects of Different Shapes
Scenario 4, the most complex of the scenarios tested, involved packing eight objects of different, irregular shapes such as gears, hooks, and rivets. The initial configuration is illustrated in Figure 24, which presents a diverse and challenging packing environment.
In this demanding scenario, the SPGD algorithm demonstrated its robust capability by converging to a significantly more compact solution compared to the traditional GD method. Although the time taken by SPGD to find the optimal solution was comparable to that of GD, the overall optimization process required more time due to the termination condition set for no improvement in the loss value over 2000 iterations. The final configurations achieved by the GD and SPGD algorithms are shown in Figure 25(a) and 25(b), reflecting the SPGD algorithm’s effectiveness in handling complex and varied object forms.
To highlight the dynamic performance of both algorithms in this scenario, Figure 26 presents the loss convergence history based on elapsed time, and the number of iterations.
These results underscore the SPGD algorithm’s capacity to adapt to and effectively manage the intricacies of packing highly irregular objects. Although the time to reach the optimal solution was similar for both algorithms, SPGD’s ability to achieve a more compact arrangement highlights its suitability for complex, real-world packing problems where shape diversity plays a critical role. The extended time required for optimization termination points to the rigorous nature of the stopping criterion, ensuring that the solution is indeed optimal before termination.
Results of these experiments are summarized in the table 5, which compares the performance of SPGD and GD in terms of best loss, and volume.
| Scenario | Method | Best Loss | Volume |
| 1 | SPGD | 74.59 | 7.12 |
| GD | 108.69 | 13.04 | |
| 2 | SPGD | 137.38 | 16.51 |
| GD | 465.53 | 58.24 | |
| 3 | SPGD | 145.59 | 17.60 |
| GD | 225.36 | 27.46 | |
| 4 | SPGD | 103.26 | 12.34 |
| GD | 123.40 | 14.76 |
This analysis highlights the superior adaptability and performance of SPGD, particularly in scenarios involving complex and non-uniform object configurations. The algorithm’s ability to effectively shuffle and perturb component sequences contributes significantly to its success in navigating the intricate landscapes presented by these diverse packing challenges.
5 Conclusion
The SPGD algorithm presents a novel integration of deterministic optimization with strategic stochastic perturbations, designed to overcome the limitations of traditional gradient descent methods in non-convex landscapes and plateaus. Through comparative analyses, SPGD has demonstrated potential advantages in complex non-convex optimization challenges, highlighting its utility across various optimization scenarios.
Looking ahead, SPGD shows promise for broader applications in diverse domains and enhancements in machine learning methodologies:
Enhancements in Machine Learning: There is potential for SPGD to significantly enhance neural network training, especially within deep learning frameworks by improving convergence rates and navigating complex parameter spaces.
Integration with Machine Learning Frameworks: SPGD has already been implemented using the PyTorch framework for the 3D component packing problem, demonstrating its adaptability to complex optimization tasks. Future work could extend this integration to machine learning projects, particularly in training neural networks, thereby potentially broadening its user base and enhancing its utility in diverse applications.
Adaptive Perturbation Strategies: Developing adaptive perturbation techniques that respond to specific characteristics of the optimization landscape could further refine SPGD’s effectiveness, making it more problem-specific.
Extension to Complex Systems: Exploring the 3D Component Packing Problem within the SPI2 framework could pave the way for handling interconnected systems with physical interactions, where topology and collision constraints add layers of complexity.
These future directions not only aim to broaden the utility of SPGD but also open new avenues for innovative research in the field of optimization.
Acknowledgments
This work was supported in part by the National Science Foundation grants CMMI-2232612, CMMI-2312175, and by the Defense Advanced Research Projects Agency (DARPA) under the “Multi-Disciplinary Optimization for Packaging (MDOP)” program, grant number FA8750-23-C-0501.